OpenStudy (anonymous):

A scientist is studying global warming. She wants to change her average temperature calculations from Celsius to Fahrenheit. The average temperature that she recorded for the Earth last year was 15 degrees Celsius. The formula for changing from Celsius to Fahrenheit is the quantity of the Celsius temperature times nine-fifths plus 32 degrees. According to the scientist’s data, what was the average temperature, in Fahrenheit, of the Earth last year?

OpenStudy (anonymous):

Simply convert Celsius to Fahrenheit using the given equation \[F = (C\cdot \frac{9}{5}) + 32\] \[F = ((15)\cdot \frac{9}{5}) + 32\] \[F = 27 + 32\] \[F = 59\:\: degrees\]

OpenStudy (anonymous):

wow thanks so much :)

OpenStudy (anonymous):

Not a problem!

OpenStudy (anonymous):

you were correct